TWO-STEP EXCITATION OF Rb AND Cs ATOMS ON He NANODROPLETS

\underline{MORITZ THEISEN}, FLORIAN LACKNER, and WOLFGANG E. ERNST, Institute of Experimental Physics, Graz University of Technology, Petersgasse 16, 8010 Graz, Austria; FRANCESCO AN-CILOTTO, Dipartimento di Fisica ‘G. Galilei’, Università di Padova, via Marzolo 8, 35131 Padova, Italy; CARLO CALLEGARI, Sincrotrone Trieste, Strada Statale 14 - km 163.5, 34149 Basovizza, Trieste, Italy.

We present the first sequential excitation of atom-doped helium nanodroplets. Rubidium atoms on the surface of a helium nanodroplet are selectively excited to the 5^{2}P_{1/2} state so as not to desorb from the droplet.^{a} From there they are excited by a laser pulse to the 5^{2}D state; a laser-induced fluorescence (LIF) spectrum is recorded by monitoring the 6^{2}P→5^{2}S_{1/2} emission. We find some difference in the LIF spectrum as compared to that of the two-photon one-color direct excitation spectrum 5^{2}D←5^{2}S_{1/2}. This indicates that the system does relax vibrationally during the lifetime of the 5^{2}P_{1/2} state. To model the LIF spectra we calculate the energy levels of the Rb atom as a function of its distance R from the center of the droplet. The Franck-Condon factors of the resulting potential energy curves agree with the experimental findings. A similar behavior has been found for cesium. New measurements predict that it also stays bound on the surface of the droplet in its 6^{2}P_{1/2} state. From there we further excited Cs monomers into their 6^{2}D state, where also the LIF spectrum is recorded by watching the 7^{2}P→6^{2}S_{1/2} emission. In the future these states can be used as a springboard to reach high-lying ^{2}S and ^{2}D states, and possibly create an artificial super-atom.^{b}

aG. Auböck, J. Nagl, C. Callegari, and W. E. Ernst, Phys. Rev. Lett. 101, 035301 (2008)

bF. Ancilotto, M. Pi, R. Mayol, M. Barranco, and K. Lehmann, J. Phys. Chem. A 111, 12695-12701 (2007)